# Distinguishing Distance Duality breaking models using electromagnetic and gravitational waves measurements

### v1

@@ 1 Introduction
The detection of gravitational wave (GW) signals has opened a new observational window into the Universe. Since the first detection by the LIGO-Virgo collaboration [1], GW have been used to investigate astrophysical phenomena [2], the nature of compact objects [3] and the behaviour of General Relativity (GR) in extreme environments [4]. The observation of GW170817 [5], originating from a binary neutron star merger with an associated electromagnetic counterpart, paved the way for the multimessenger use of gravitational waves, enabling cosmologists to utilize them as an independent probe for a measurement of the Hubble constant, $H_{0}$ [6], opening a new avenue for the use of GW for cosmological investigation. The combination of gravitational wave data with electromagnetic observations in GW170817 significantly improved the precision of the $H_{0}$ estimate — by a factor of 2–3 — compared to analyses based on gravitational waves alone [7, 8]. In addition to this, the improved detection capabilities of future GW observatories, such as the Einstein Telescope [9] or Cosmic Explorer [10], will open new frontiers in the study of the cosmos.
A key application of GW in cosmology is their role as standard sirens, which allows to directly measure the luminosity distance ($d_{L}$), without the need for a calibration as is necessary for electromagnetic standard candles such as type Ia Supernovae (SNe). The detection of GW signals and (when possible) of their electromagnetic counterparts enables independent constraints on cosmological parameters and, consequently, permits to investigate the nature of dark energy as well as fundamental questions about the gravitational interaction [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Besides obtaining independent constraints on the standard cosmological model $\Lambda$CDM, able to describe the Universe’s history in terms of cold dark matter (CDM) and the cosmological constant ($\Lambda$), with gravity described by GR, this new observational probe is also extremely promising to test the fundamental assumptions at the basis of this description. Among the observational consequences of such fundamental assumptions, one that has gathered significant interest is the so-called Distance Duality Relation (DDR), or Etherington’s relation [23], i.e. the relation between $d_{L}$ and angular diameter distance ($d_{A}$), which spawns directly from the assumptions of the standard cosmological model, and that is not violated in simple extensions.
@@
Commonly, violations of the DDR have been explored in the electromagnetic (EM) sector, where they could signal the necessity of new physics, such as the non-conservation of the photon number, deviations from a metric theory of gravity, or unusual photon propagation effects [24]. Several works have investigated how current and upcoming measurements of luminosity and angular distances can be combined together to obtain tight constraints on DDR deviations [25, 26, 27, 28], while others have focused on the search of stand-alone probes for such a non-standard behaviour, such as strongly lensed systems [29, 30]. Furthermore, violations of the DDR have been explored also in the context of the $H_{0}$ tension [31, 32, 33], investigating how the latter could be an observational consequence of a failure of the assumptions leading to the DDR [34, 35].
@@
Even though most of the investigations done in the DDR context have focused on EM probes, the expected improvement in GW measurements has prompted cosmologists to enquire how these can be used to further test such a fundamental relation. In particular, recent results have highlighted how GW could be used to break the degeneracy between violations of the DDR and the impact of modified gravity theories on observations [36], while it has also been shown that combining measurements of EM and GW distances could lead to tight and model-independent tests of the DDR [37].
@@
In this work, we focus on assessing the improvement in constraining power on DDR violations when GW, in the form of bright sirens, are used alongside EM probes to bound parameterised deviations from standard behaviour. We show, however, how this implies the assumption that the mechanism responsible for possible violations affects the gravitational and luminosity distances in the same way. As this is not necessarily the case, with models breaking the DDR differently in the EM and GW sectors [38], a key result of this work is to show the role of GW in distinguishing between these possibilities, offering a way to identify whether a hypothetical violation of the DDR arises from electromagnetic or gravitational origins. Given the sensitivity required for this kind of investigation, we restrict our analysis to upcoming surveys in both the EM and GW sectors. In particular, we rely on the expected sensitivity of the Einstein Telescope [39] for the measurement of GW distances, of SKAO for estimates of the angular diameter distance through baryon acoustic oscillations (BAO) measurements [40], and on the the expected SN observations of the Vera C. Rubin observatory [41], which will give information on the EM luminosity distance.
@@ 2 Distance duality relation and possible violations
In any cosmological model based on a metric theory of gravity, the Etherington relation, or distance duality relation, implies that distance measures are unique [23]. Specifically, at any redshift $z$, a well-defined relation connects the angular diameter distance $d_{A}(z)$, which characterizes the apparent size of a distant source, to the luminosity distance $d_{L}(z)$, which is determined by the flux received from that source. In the framework of GR, the number of particles emitted from the source is conserved, and they travel on null geodesics (i.e., they are massless). In this case, the DDR implies
@@
While DDR violations could be associated with the failure of multiple assumptions at the basis of standard cosmology, the violation of different assumptions would lead to distinct observational signatures. In particular, it is important to stress that the luminosity distance can be connected to electromagnetic observables (e.g., type Ia SNe), but also to measurements of GW, which would not rely on the observations of photons. For such a reason, we distinguish here between two types of luminosity distances: electromagnetic $d_{\rm L}^{\rm EM}(z)$ and gravitational $d_{\rm L}^{\rm GW}(z)$. While these two quantities coincide in standard cosmology, a violation of DDR would in general affect the two distances differently, depending on the mechanism leading to the breakdown of cosmological assumptions. As an example, in models where photons can decay into another particle that remains unobserved, there is an effective violation of photon number conservation, which leads to a departure of $\eta^{\rm EM}(z)$ from unity [38]. At the same time, the propagation of GW is not affected by such a mechanism and, therefore, $\eta^{\rm GW}=1$. On the contrary, mechanisms that leads to violation of the DDR through modifications of the space-time geometry, e.g. in torsion extensions of GR [42], can impact the evolution of both EM and GW distances, thus leading to a different signature in the GW sector with respect to mechanisms that only affect photons.
@@ 2.1 Phenomenological deviations from the DDR
The deviation from the standard DDR is encoded in $\epsilon(z)$, which is commonly taken as a constant. The constraints obtained with current data are of the order of the order of $10^{-1}$ in the $2\sigma$ range [25].
@@
In order to obtain constraints on these functions, a common approach is to parameterize them. In the easiest approach, the function $\epsilon(z)$ is assumed to be a constant, with its value $\epsilon_{0}$ being a free parameter to be constrained with observational data [24, 38, 25]. However, with the increase in precision of the available data, different approaches have been used, trying to also detect possible redshift trends for $\epsilon(z)$ or $\eta(z)$, e.g. by binning in redshift [25] or applying machine learning reconstruction methods such as Gaussian Processes or Genetic Algorithms [36, 29]. While this phenomenological definition applies in general, it is important to explore the theoretical models leading to violations of the DDR, as their impact on cosmological observables will be different.
@@ 2.1.1 Padé parametrization
It has been argued that the trend of DDR violations produced by physical models should be monotonic in redshift [43]. While a monotonic trend can be reproduced by the parametrization we presented above, such a behaviour can lead to issues when modelling theoretical predictions for high redshift observables, which would require studying physical mechanisms in such a modified scenario. It is therefore common to work with parametrizations that include a cut-off of the DDR violation, neglecting the impact of DDR violations at high redshift. In order to avoid an artificially sharp cut-off, it is possible to rely on approaches that naturally lead the modifications introduced to vanish at high redshift.
A robust approach leading to such an effect involves the use of rational polynomials [44, 45]. A notable example in this respect is offered by Padé approximants, which have been applied to cosmographic reconstructions of the Hubble expansion rate to heal the convergence issues inherent in Taylor series expansions of the luminosity distance [46, 47], as they offer both stability at large cosmological distances and improved convergence properties [48, 49].
The Padé approximant method has found broad applications across various contexts, ranging from kinematic analyses in standard cosmology to investigations within modified gravity frameworks [50, 51, 52, 53]. Specifically, one can construct the $(n,m)$ Padé approximant of a function $f(z)$ as
@@
[…]
@@
Motivated by the optimal convergence properties and statistical performances of the (2,1) Padé polynomial as demonstrated in [47], we consider the following parametrization:
@@
In 1(a), we show the behavior of $\eta(z)$ when modeled using a Padé approximation compared to a simpler expression given in Eq. (2.7). The inset in the figure highlights the differences between the two approaches within the redshift range relevant to our analysis. Additionally, in 1(b), we present the relative difference between the two parametrizations.
Figure 1: (a) : Comparison of the behaviour of $\eta(z)$ using the Padé expression (blue) and the constant $\epsilon$ expression. The inset shows a zoomed-in view of the redshift range relevant for this work. The horizontal grey line marks $\eta=1$, and the vertical dashed grey line indicates the redshift of the Last Scattering Surface $z_{*}$. (b) : Relative difference between $\eta(z)$ computed using a constant $\epsilon$ model (Const) and the Padé expression (Padé) across the redshift range $0<z<z_{*}$.
@@ 3 Cosmological data
We therefore use two different observables to obtain measurements related to luminosity distances. On the one hand, gravitational waves observations will provide measurement of $d_{\rm L}^{\rm GW}(z)$, i.e. the distance that can be measured from the merging of compact binaries, while SN-Ia observations are related to $d_{\rm L}^{\rm EM}(z)$, through their use as standardisable candles. While these two observables are in principle enough to obtain constraints on the violation of DDR for the two sectors, the DDR parameters are very degenerate with other standard cosmological parameters, e.g. with the energy density of matter $\Omega_{\rm m}$ [25]. For such a reason, in order to obtain tight constraints, we include in our set of observables also BAO observations. The latter provide measurements of quantities that are related to $d_{A}(z)$ and $H(z)$, which are not affected by violations of the DDR. For this reason, BAO can act as an anchor for the constraints obtained with GW and SN, breaking the degeneracies with cosmological parameters. In this work, we utilize these three observables, with a particular emphasis on future datasets. In subsection 3.1, subsection 3.2, subsection 3.3 we enter into the details of each observable, in particular on how distance measurements can be extracted from observations, and we explain the methodology that we follow to simulate data from upcoming surveys.
@@ 3.1 Baryon acoustic oscillations
BAO observations measure the quantities $D_{M}/r_{d}$, $D_{H}/r_{d}$, $D_{V}/r_{d}$, where the distances in a flat Universe are defined as [54]:
@@
As said before, the BAO measurements are not sensitive to violation in the DDR relation but they help in breaking the degeneracy between the DDR parameter $\epsilon$ and $\Omega_{m}$, as shown in previous works [25] and in Figure 2
@@
In this work, we explore the constraining power of future BAO measurements, specifically those expected from the SKA Observatory (SKAO) [55]. In particular, we simulate the BAO measurements assuming the optimistic configuration, labelled as SKA2 in [40], where the 21-cm spectral line will be probed to extract BAO information in the redshift range $z\in\left[0.2,2\right]$, divided in $18$ bins of width $\Delta z=0.1$.
It must be pointed out that BAO measurements do not provide information directly on the distances of Eq. (3.3), but rather on their ratio with the sound horizon $r_{d}$. As it can be seen from Eq. (3.1) and Eq. (3.2), disentangling such a degeneracy would require information on the baryonic energy density $\rho_{b}$, which cannot be provided by BAO measurements. Indeed, the use of external priors, such as those coming from Big Bang Nucleosynthesis (BBN) [56], allows to probe directly the BAO distances, but we choose here to avoid the use of external prior as it would require us to model the impact of a DDR violation on early Universe physics. For such a reason, we simulate a dataset for $D_{M}/r_{d}$ and $D_{H}/r_{d}$, propagating to these quantities the relative errors on $d_{A}$ and $H$ presented in Figure 4 of [57].

### v2

@@ 1 Introduction
The detection of gravitational wave (GW) signals has opened a new observational window into the Universe. Since the first detection by the LIGO-Virgo collaboration [LIGOScientific:2016aoc], GW have been used to investigate astrophysical phenomena [Abbott_2023], the nature of compact objects [Abbott_properties] and the behaviour of General Relativity (GR) in extreme environments [Abbott_2016]. The observation of GW170817 [LIGOScientific:2017vwq], originating from a binary neutron star merger with an associated electromagnetic counterpart, paved the way for the use of gravitational waves in the context of multimessenger cosmology. This allowed cosmologists to exploit GW observations as an independent probe for a measurement of the Hubble constant, $H_{0}$ [LIGOScientific:2017adf], opening a new avenue for cosmological investigation. Indeed, the combination of gravitational wave data with electromagnetic observations in GW170817 significantly improved the precision of the $H_{0}$ estimate — by a factor of 2–3 — compared to analyses based solely on gravitational waves [Hotokezaka2018, Gianfagna_2024]. In addition to this, the improved detection capabilities of future GW observatories, such as the Einstein Telescope [Abac:2025saz] or Cosmic Explorer [Reitze:2019iox], will open new frontiers in the study of the cosmos.
[…]
@@
Commonly, violations of the DDR have been explored in the electromagnetic (EM) sector, where they could signal the necessity of new physics, such as the non-conservation of the photon number, deviations from a metric theory of gravity, or unusual photon propagation effects [Avgoustidis2009]. Several works have investigated how current and upcoming measurements of luminosity and angular distances can be combined together to obtain tight constraints on DDR deviations [EUCLID:2020syl, Favale:2024sdq, Keil:2025ysb, Alfano:2025gie], while others have focused on the search of stand-alone probes for such a non-standard behaviour, such as strongly lensed systems [Renzi:2020bvl, Tang:2024zkc]. Furthermore, violations of the DDR have been explored also in the context of the $H_{0}$ tension [Perivolaropoulos:2021jda, Abdalla:2022yfr, DiValentino:2025sru], investigating how the latter could be an observational consequence of a departure from the DDR [Renzi:2021xii, Teixeira:2025czm]. Violations of the DDR have also been investigated in the context of addressing potential statistical correlations among cosmographic parameters [Jesus_2025].
@@
Even though most of the investigations done in the context of DDR have focused on EM probes, the expected improvement in GW measurements has prompted cosmologists to enquire how these can be used to further test such a fundamental relation [yang2019constraintscosmicdistanceduality]. In particular, recent results have highlighted how GW could be used to break the degeneracy between violations of the DDR and the impact of modified gravity theories on observations [Hogg:2020ktc], while it has also been shown that combining measurements of EM and GW distances could lead to tight and model-independent tests of the DDR [Matos:2023jkn]. Finally, other studies have shown that GW events from current and upcoming detectors can serve as a model-independent test of General Relativity by probing the relationship between the luminosity distance in the gravitational and EM sectors [afroz2024prospectprecisioncosmologytesting, afroz2024modelindependentprecisiontestgeneral, afroz2024modelindependentprecisiontestgeneral2].
@@
In this work, we focus on assessing the improvement in constraining power on DDR violations when GW, in the form of bright sirens, are used alongside EM probes to bound parameterised deviations from standard behaviour. We show, however, how this implies the assumption that the mechanism responsible for possible violations affects the gravitational and luminosity distances in the same way. As this is not necessarily the case, with models breaking the DDR differently in the EM and GW sectors [Avgoustidis2010], a key result of this work is to show the role of GW in distinguishing between these possibilities, offering a way to identify whether a hypothetical violation of the DDR arises from electromagnetic or gravitational origins. Given the sensitivity required for this kind of investigation, we restrict our analysis to upcoming surveys in both the EM and GW sectors. In particular, we rely on the expected sensitivity of the Einstein Telescope [abac2025scienceeinsteintelescope] for the measurement of GW distances, of SKAO for estimates of the angular diameter distance through baryon acoustic oscillations (BAO) measurements [SKAO_gen], and on the expected SN observations of the Vera C. Rubin observatory [Bianco2022], which will give information on the EM luminosity distance.
@@ 2 Distance duality relation and possible violations
In any cosmological model based on a metric theory of gravity, the Etherington relation, or distance duality relation, implies that distance measures are unique [Etherington]. Specifically, at any redshift $z$, a well-defined relation connects the angular diameter distance $d_{A}(z)$, which characterizes the apparent size of a distant source, to the luminosity distance $d_{L}(z)$, which is determined by the flux received from that source. In the framework of GR, the number of particles emitted from the source is conserved, and they travel on null geodesics (i.e., they are massless). In this case, the DDR implies
@@
While DDR violations could be associated with the failure of multiple assumptions at the basis of standard cosmology, the violation of different assumptions would lead to distinct observational signatures. In particular, it is important to stress that the luminosity distance can be connected to electromagnetic observables (e.g., type Ia SNe), but also to measurements of GW, which would not rely on the observations of photons. For such a reason, we distinguish here between two types of luminosity distances: electromagnetic $d_{\rm L}^{\rm EM}(z)$ and gravitational $d_{\rm L}^{\rm GW}(z)$. While these two quantities coincide in standard cosmology, a violation of DDR would in general affect the two distances differently, depending on the mechanism leading to the breakdown of cosmological assumptions. As an example, in models where photons can decay into another particle that remains unobserved, there is an effective violation of photon number conservation, which leads to a departure of $\eta^{\rm EM}(z)$ from unity [Avgoustidis2010]. At the same time, the propagation of GW is not affected by such a mechanism and, therefore, $\eta^{\rm GW}=1$. On the other hand, mechanisms that leads to violation of the DDR through modifications of the space-time geometry, e.g. in torsion extensions of GR [Bolejko:2020nbw, DAgostino:2025kme], can impact the evolution of both EM and GW distances, thus leading to a different signature in the GW sector with respect to mechanisms that only affect photons.
More generally, several classes of modified gravity theories predict a deviation of the GW luminosity distance from the standard form. For instance, scalar-tensor theories (including Horndeski and beyond-Horndeski models) can alter GW propagation through a time-varying Planck mass or additional friction terms, leading to a redshift-dependent modification of $d_{L}^{\rm GW}(z)$ [Ezquiaga:2017ekz, Belgacem:2018lbp, Nishizawa:2019rra]. Similarly, in massive gravity and bimetric theories, a modified dispersion relation for GWs or coupling to an auxiliary metric can lead to observable effects in the GW sector [deRham:2010kj, Hassan:2011zd, Gumrukcuoglu:2012wt, deRham:2014zqa]. These theoretical frameworks provide well-motivated mechanisms in which the breaking of the DDR can be tested with multi-messenger observations. We therefore emphasize that the comparison between $d_{L}^{\rm{EM}}(z)$ and $d_{L}^{\rm{GW}}(z)$ offers a powerful diagnostic tool to discriminate among the physical origins of a possible DDR violation, distinguishing, e.g., between photon-specific effects and modifications to the gravitational sector.
@@ 2.1 Phenomenological deviations from the DDR
The deviation from the standard DDR is encoded in $\epsilon(z)$, which is commonly taken as a constant. The constraints obtained with current data are of the order of the order of $10^{-1}$ in the $2\sigma$ range [EUCLID:2020syl].
@@
In order to obtain constraints on these functions, a common approach is to parameterize them. In the easiest approach, the function $\epsilon(z)$ is assumed to be a constant, with its value $\epsilon_{0}$ being a free parameter to be constrained with observational data [Avgoustidis2009, Avgoustidis2010, EUCLID:2020syl]. However, with the increase in precision of the available data, different approaches have been used, trying to also detect possible redshift trends for $\epsilon(z)$ or $\eta(z)$, e.g. by binning in redshift [EUCLID:2020syl] or applying machine learning reconstruction methods such as Gaussian Processes or Genetic Algorithms [Hogg:2020ktc, Renzi:2020bvl]. While this phenomenological definition applies in general, it is important to explore the theoretical models leading to violations of the DDR, as their impact on cosmological observables will be different.
@@ 2.1.1 Padé parametrization
It has been argued that the trend of DDR violations produced by physical models should be monotonic in redshift [Avgoustidis:2009ai]. While a monotonic trend can be reproduced by the parametrization we presented above, such a behaviour can lead to issues when modelling theoretical predictions for high redshift observables, which would require studying physical mechanisms in such a modified scenario. It is therefore common to work with parametrizations that include a cut-off of the DDR violation, neglecting the impact of DDR violations at high redshift. In order to avoid an artificially sharp cut-off, it is possible to rely on approaches that naturally lead the modifications introduced to vanish at high redshift.
A robust approach leading to such an effect involves the use of rational polynomials [Wei:2013jya, Capozziello:2017nbu]. A notable example in this respect is offered by Padé approximants, which have been applied to cosmographic reconstructions of the Hubble expansion rate to heal the convergence issues inherent in Taylor series expansions of the luminosity distance [Capozziello:2022jbw, Capozziello:2020ctn], as they offer both stability at large cosmological distances and improved convergence properties [Aviles:2014rma, DAgostino:2023cgx].
The Padé approximant method has found broad applications across various contexts, ranging from kinematic analyses in standard cosmology to investigations within modified gravity frameworks [Capozziello:2017ddd, Capozziello:2019cav, Capozziello:2022wgl, Capozziello:2023ccw]. Specifically, one can construct the $(n,m)$ Padé approximant of a function $f(z)$ as
@@
$\left\{\begin{aligned} &a_{i}=\sum_{k=0}^{i}b_{i-k}\ c_{k}\ ,\\ &\sum_{j=1}^{m}b_{j}\ c_{n+k+j}=-b_{0}\ c_{n+k}\ ,\quad k=1,\ldots,m\end{aligned}\right$. (2.9)
@@
Motivated by the optimal convergence properties and statistical performances of the (2,1) Padé polynomial as demonstrated in [Capozziello:2020ctn], we consider the following parametrization:
@@
In Figure 1, we show the behavior of $\eta(z)$ when modeled using a Padé approximation compared to a simpler expression given in Eq. (2.7). The inset in the figure highlights the differences between the two approaches within the redshift range relevant to our analysis.
Figure 1: Comparison of the behaviour of $\eta(z)$ using the Padé expression (blue) and the constant $\epsilon$ expression. The inset shows a zoomed-in view of the redshift range relevant for this work. The horizontal grey line marks $\eta=1$, and the vertical dashed grey line indicates the redshift of the Last Scattering Surface $z_{*}$.
@@ 3 Cosmological data
We therefore use two different observables to obtain measurements related to luminosity distances. On the one hand, gravitational waves observations will provide measurement of $d_{\rm L}^{\rm GW}(z)$, i.e. the distance that can be measured from the merging of compact binaries, while SN-Ia observations are related to $d_{\rm L}^{\rm EM}(z)$, through their use as standardisable candles. While these two observables are in principle enough to obtain constraints on the violation of DDR for the two sectors, the DDR parameters are very degenerate with other standard cosmological parameters, e.g. with the energy density of matter $\Omega_{\rm m}$ [EUCLID:2020syl]. For such a reason, in order to obtain tight constraints, we include in our set of observables also BAO observations. The latter provide measurements of quantities that are related to $d_{A}(z)$ and $H(z)$, which are not affected by violations of the DDR. For this reason, BAO can act as an anchor for the constraints obtained with GW and SN, breaking the degeneracies with cosmological parameters. In this work, we utilize these three observables, with a particular emphasis on future datasets. In subsection 3.1, subsection 3.2, subsection 3.3 we enter into the details of each observable, in particular on how distance measurements can be extracted from observations, and we explain the methodology that we follow to simulate data from upcoming surveys.
@@ 3.1 Baryon acoustic oscillations
BAO observations measure the quantities $D_{M}/r_{d}$, $D_{H}/r_{d}$, $D_{V}/r_{d}$, where the distances in a flat Universe are defined as [DESI:2024mwx]:
@@
As said before, the BAO measurements are not sensitive to violation in the DDR relation but they help in breaking the degeneracy between the DDR parameter $\epsilon$ and $\Omega_{m}$, as shown in previous works [EUCLID:2020syl] and in Figure 2
@@
In this work, we explore the constraining power of future BAO measurements, specifically those expected from the SKA Observatory (SKAO) [SKAO]. In particular, we simulate the BAO measurements assuming the optimistic configuration, labelled as SKA2 in [SKAO_gen], where the 21-cm spectral line will be probed to extract BAO information in the redshift range $z\in\left[0.2,2\right]$, divided in $18$ bins of width $\Delta z=0.1$.
It must be pointed out that BAO measurements do not provide information directly on the distances of Eq. (3.3), but rather on their ratio with the sound horizon $r_{d}$. As it can be seen from Eq. (3.1) and Eq. (3.2), disentangling such a degeneracy would require information on the baryonic energy density $\rho_{b}$, which cannot be provided by BAO measurements. Indeed, the use of external priors, such as those coming from Big Bang Nucleosynthesis (BBN) [BBN], allows to probe directly the BAO distances, but we choose here to avoid the use of external prior to be consistent with the choice of neglecting the SH0ES calibration for the SNe, as we will discuss in subsection 3.2. While an early-time DDR violation would impact both the BBN process and the BAO scale, we neglect these effects here as they lie outside the scope of this work. For such a reason, we simulate a dataset for $D_{M}/r_{d}$ and $D_{H}/r_{d}$, propagating to these quantities the relative errors on $d_{A}$ and $H$ presented in Figure 4 of [Bull:2015lja].